# Latent Multi-Head Attention for Small Language Models

### v1

@@ Appendix B Extended GPT-4 Evaluation Results
MLA+RoPE maintains remarkably stable quality through moderate compression levels. At $r=d/2$, the model retains 97% of its original overall score, demonstrating the effectiveness of learned compression with positional information. Even at $r=d/4$, the model achieves 92% of baseline performance, remaining fully functional for practical applications.
@@
Across all configurations, narrative consistency shows the steepest degradation with compression, serving as an early indicator of model breakdown. For MLA without RoPE, consistency drops from 4.8 to 2.0 by $r=d/16$, while grammar maintains relative stability at 3.5. This aligns with our theoretical understanding that maintaining long-range dependencies requires sufficient attention capacity.
@@
Both architectures exhibit complete failure at $r=d/32$, with scores approaching the theoretical minimum. The generated text at this compression level consists of repetitive tokens, grammatically broken fragments, and no discernible narrative structure. The slightly higher scores for MLA+RoPE (2.3 vs 1.6) reflect occasional coherent phrases rather than meaningful generation capability.
@@
The performance gap between MLA and MLA+RoPE widens dramatically at higher compression ratios. While both start with a  2 point difference at $r=d$, this gap expands to  1.4 points at $r=d/16$, demonstrating that rotary embeddings provide crucial structural information that partially compensates for reduced attention capacity.
Appendix C Qualitative Examples
We provide representative story completions from each model variant (12L-1024d) to illustrate the quality differences discussed in Section 3.4:

### v2

@@ Appendix B Appendix B: Extended GPT-4 Evaluation Results
MLA+RoPE maintains remarkably stable quality through moderate compression levels. At $r=d/2$, the model retains 97% of its original overall score, demonstrating the effectiveness of the learned compression with positional information. Even at $r=d/4$, the model achieves 92% of baseline performance, remaining fully functional for practical applications.
@@
Across all configurations, narrative consistency shows the steepest degradation with compression, serving as an early indicator of model breakdown. For MLA without RoPE, consistency drops from 4.8 to 2.0 by $r=d/16$, while the grammar maintains relative stability at 3.5. This aligns with our theoretical understanding that maintaining long-range dependencies requires sufficient attention capacity.
@@
Both architectures exhibit complete failure at $r=d/32$, with scores approaching the theoretical minimum. The text generated at this compression level consists of repetitive tokens, grammatically broken fragments, and no discernible narrative structure. The slightly higher scores for MLA + RoPE (2.3 vs. 1.6) reflect occasional coherent phrases rather than meaningful generation capability.
@@
The performance gap between MLA and MLA+RoPE widens dramatically at higher compression ratios. Although both start with a  2 point difference at $r=d$, this gap expands to  1.4 points at $r=d/16$, demonstrating that rotary embeddings provide crucial structural information that partially compensates for reduced attention capacity.
Appendix C Appendix C: Qualitative Examples
We provide representative story completions for each model variant (12L-1024d) to illustrate the quality differences discussed in Section 3.4: